Alden Co.’s monthly unit sales and total cost data for its operating activities of the past year follow. Management wants to use these data to predict future fixed and variable costs.

Month Units Sold Total Cost Month Units Sold Total Cost
1 320,500 $ 158,000 7 359,500 $ 274,364
2 165,500 101,750 8 270,500 152,250
3 265,500 206,100 9 75,900 64,500
4 205,500 100,500 10 150,500 131,125
5 290,500 202,000 11 94,500 94,500
6 190,500 112,500 12 100,500 81,150

Please help me with this 2 part question

1. Estimate both the variable costs per unit and the total monthly fixed costs using the high-low method. (Do not round intermediate calculations.)

2. Predict future total costs when sales volume is (a) 389,000 units and (b) 429,000 units.
Answer #1

Answer

  • High Low Method working

Months

Units

Cost

High Level

7

                        359,500

$                274,364.00

Low Level

9

                          75,900

$                   64,500.00

Difference

                        283,600

$                209,864.00

A

Difference in Cost

$                209,864.00

B

Difference in units

                          283,600

C = A/B

Variable cost per unit

$                           0.740

Working

High Level

Low Level

A

Total Cost

$              274,364.00

$                   64,500.00

B

Total Units

359500

75900

C

Variable cost per unit

$                           0.74

$                             0.74

D = B x C

Total Variable cost

$              266,030.00

$                   56,166.00

E = A - D

Total Fixed Cost

$                   8,334.00

$                     8,334.00

  • Requirement 1

Variable cost per unit = $ 0.74

Monthly Fixed Cost = $ 8,334

  • Requirement 2

[a] Cost for 389000 units = $ 8334 + ($0.74 x 389000) = $ 296,194
[b] Cost for 429000 units = $ 8334 + ($0.74 x 429000) = $ 325,794